| Operation Summary for March 26, 2012 - March 28, 2012 |
|
Monday March 26 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, SeaQuest experiments, and the FTBFT-1017 experiment taking beam. During the day shift, Operators prepared SeaQuest for access at 10:01 AM. Operations reported at 11:24 AM that the P1 transfer line suffered a vacuum burst; Operations halted the Switchyard (SY) beam for about 13 minutes while the P1 transfer line vacuum recovered. Three Linac quadrupole magnet power supplies (104, 108, and 114) tripped off at 12:14 PM; they reset. At 12:18 PM, a SY SEPTA (S:FSEPV) tripped off on a no control power indication; Power Supply personnel investigated and replaced a five-volt power supply. Operations established beam to MTA at 1:22 PM. SeaQuest came out of access and resumed taking beam at 3:41 PM. Booster personnel conducted a study during this shift and T-1017 made an access during this shift. During the evening shift, at 4:04 PM, six Linac corrector magnet (L:HT1N, L:VTIN, L:HT3N, L:VT3N, L:HT4N, and L:VT4N) tripped off; they all reset. The Booster study ended at 5:23 PM. Operations disabled beam to MTA at 5:30 PM. At 6:11 PM, the NuMI beam permit tripped off for the first of two times in 26 minutes due to a power supply’s output (E:V118) drifting low; everything reset. Tuesday March 27 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1017. During the midnight shift, at 12:36 AM, the MiniBooNE beam permit tripped off due to a MI-12 enclosure airflow interlock for the first of two times in one minute; Operations contacted a MiniBooNE person who had an Operator make adjustments. At 2:24 AM, the NuMI beam permit tripped off due to V118; everything reset. At 6 AM, the MI beam permit tripped off due to a bad beam position and NuMI tripped off due to an incomplete RF gap; everything reset. During the day shift, Muon Ring personnel began a muon target yield study using the antiproton target at 9:47 AM. At 11:04 AM, the MiniBooNE beam permit tripped off for the first of seven times in 38 minutes due to an MI-12 airflow alarm; it reset. A MI study ended at 1:31 PM. At 2:30 PM, the NuMI beam permit tripped off for the first of two times in about an hour due to MI RF station trips (MIRF7); everything reset. SeaQuest made accesses during this shift. During the evening shift, the Muon Ring study ended at 4:46 PM. At 5:05 PM, the NuMI beam permit tripped off for the first of three times due to V118; it reset. At 5:19 PM, MIRF7 tripped off for the first of 10 times in six hours; it reset. At 10:02 PM, the H- Source shut down when its anode power supply tripped off; Linac personnel came in and had the source on by 10:46 PM. Wednesday March 28 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1017. During the midnight shift, at 12:36 AM, MIRF7 tripped off for the first of five times in about two hours; Operators reset the station and after the last trip adjusted the station’s anode program. At 1:40 AM, the NuMI beam permit tripped off for the first of four times in about four hours due to V118; it reset. A MI kicker (I:KPS6N) tripped off at 2:30 AM; it reset. T-1017 reported at 3:54 AM that they had completed their current run at Fermilab. Operations Coordinator (OpsCo) Plans for Today The OpsCo agrees with MI and Muon Ring study plans for today. |
|
More Information Comments and Suggestions |
| last modified 03/28/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |